# RevisEval: Improving LLM-as-a-Judge via Response-Adapted References

### v1

@@ Abstract
With significant efforts in recent studies, LLM-as-a-Judge has become a cost-effective alternative to human evaluation for assessing the text generation quality in a wide range of tasks. However, there still remains a reliability gap between LLM-as-a-Judge and human evaluation. One important reason is the lack of guided oracles in the evaluation process. Motivated by the role of reference pervasively used in classic text evaluation, we introduce RevisEval, a novel text generation evaluation paradigm via the response-adapted references. RevisEval is driven by the key observation that an ideal reference should maintain the necessary relevance to the response to be evaluated. Specifically, RevisEval leverages the text revision capabilities of large language models (LLMs) to adaptively revise the response, then treat the revised text as the reference (response-adapted reference) for the subsequent evaluation. Extensive experiments demonstrate that RevisEval outperforms traditional reference-free and reference-based evaluation paradigms that use LLM-as-a-Judge across NLG tasks and open-ended instruction-following tasks. More importantly, our response-adapted references can further boost the classical text metrics, e.g., BLEU and BERTScore, compared to traditional references and even rival the LLM-as-a-Judge. A detailed analysis is also conducted to confirm RevisEval’s effectiveness in bias reduction, the impact of inference cost, and reference relevance.
@@ 1 Introduction
As the large language model (LLM) already exhibits strong alignment with humans (Gilardi et al., 2023; OpenAI et al., 2024), LLM-as-a-Judge (Chang et al., 2024; Li et al., 2024b; Gao et al., 2024), aka. LLM-evaluator, has emerged as a viable alternative to human evaluation in assessing text generation quality. Given the task instruction and the corresponding model-generated responses, LLMs are prompted to predict preferences or scores for these responses. Despite considerable efforts have been made, such as chain-of-thought (Zheng et al., 2023), specialized rubrics (Liu et al., 2023), and extensive evaluation-specific training datasets (Li et al., 2024a; Wang et al., 2024c; b), human evaluation remains the gold standard in text quality assessment (Zeng et al., 2024) and LLM-as-a-Judge struggles with particular biases (Huang et al., 2024) and being vulnerable to the misleading context (Dubois et al., 2024; Chen et al., 2024). One important reason is the lack of an oracle to direct the evaluation process. Fortunately, classical text evaluation metrics, like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), offer valuable prompts in mitigating such a gap: given an appropriate reference, i.e., the ground-truth answer to the task, calculating the similarity between the references and the model-generated responses can achieve a satisfactory correlation with human evaluations. Furthermore, several studies highlight the reference could prevent being overly sensitive to semantic deficiency  (Sheng et al., 2024) and overcoming bias (Deutsch et al., 2022) in certain cases.
@@
Motivated by the above findings, we deem that an effective reference should maintain high quality while ensuring relevance to the response, which led us to consider that revising the response adaptively could be a good candidate for this reference (Guo et al., 2024). Therefore, we propose a novel evaluation paradigm Revise-and-Evaluation, abbreviated as RevisEval. Specifically, given the (instruction, response) pair, RevisEval first revise the response using the instruction and evaluation rubric, resulting in the response-adapted reference. RevisEval further leverages this generated response-adapted reference to guide final evaluation (e.g., scoring or pairwise comparison). By revising the original response, we can ensure that the generated reference is both high-quality and closely relevant to the original content. The comparison between the original and revised responses offers valuable insights for evaluation. Orthogonal to previous work that only focuses on the discrimination abilities of LLMs, RevisEval stands out by fully utilizing the generative potential by revision.
We conduct comprehensive experiments to validate the effectiveness of our proposed RevisEval. Using both proprietary and open-source LLMs, RevisEval consistently achieves better evaluation performance compared to reference-free and reference-based evaluation paradigms in both NLG tasks and open-ended instruction-following tasks. Moreover, we seek to verify the effectiveness of the response-adapted references in the classic metrics, e.g., BERT and BERTScore, showing that each metric exceeds itself by up to 3%-10% accuracy compared to using direct response as references. We then combine LLM-as-a-reviser with multiple classic metrics and find that it outperforms LLM-as-a-Judge (reference-free setting) by over 1.5% on average when using weak LLMs and is comparable when using GPT-4. Finally, we analyze how our paradigm achieves overall superiority. In reducing verbosity and positional bias, our approach offers clear advantages in adversarially designed LLMBar and swapping position testing. Merely increasing inference cost of reference-free evaluation still lags behind RevisEval, demonstrating our method’s efficiency does not rely on naively accumulating cost. Meanwhile, we validate the relationship between reference relevance and efficiency using response-adapted references.
@@ 2.1 Evaluation of Large Language Models
Instruction-tuned LLMs (Ouyang et al., 2022; Dubey et al., 2024; Team et al., 2023) have revolutionized the field of natural language processing (NLP) due to their ability to handle a wide range of language-related tasks. Unlike traditional NLP tasks, such as Machine Translation and Summarization, which can be evaluated using N-gram-based metrics like BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) by comparing responses with reference texts, LLMs excel at open-ended language generation tasks (e.g., story generation and open-ended instruction-following generation), where no single reference response exists. Consequently, several studies embrace the potential of LLM-as-a-Judge and shift toward reference-free metrics, advocating the abandonment of conventional reference-based evaluation methods (Sheng et al., 2024; Chen et al., 2023). In this paper, we re-visit the significance of reference in LLM evaluation. Furthermore, to address the challenge of the absence of a single standard answer in certain evaluation tasks, we propose leveraging LLMs to generate response-adapted intermediate references, thereby improving the evaluation performance of both traditional metrics and LLM-as-a-Judge.
@@ 3 Methodology
Figure 2: Illustration of our proposed RevisEval. Given an instance ($x,y,a$), we use RevisEval to assess $y$ in rubric $a$. In RevisEval, (i) reviser generates a response-adapted reference $r^{\star}$ by revising the $y$ to enhance the (ii) following LLM-as-a-Judge, even classic text metrics.
@@ 4 Experiments
In this section, we first present the comprehensive experimental settings in Sec. 4.1 and evaluate the RevisEval in LLM-as-a-Judge across various tasks in Sec. 4.2; we then verify the effectiveness of classic text evaluation metrics when using response-adapted references in Sec. 4.3; building on above findings, we compare two evaluation paradigms, combining LLM-as-a-reviser with multiple metrics and LLM-as-a-Judge, when using weak LLM in Sec. 4.4; finally, we conduct detailed comparative analysis of RevisEval in Sec. 4.5, such as bias reduction, inference cost impact and reference relevance.
@@ 4.1 Evaluation Settings
We evaluate our approach on multiple classic NLG benchmarks by measuring the correlation between the evaluators/metrics and human annotations in a scoring rating task. We follow the experimental setting of Jiang et al. (2024a) and select four representative NLG tasks: Data-to-Text, Machine Translation, Text Summarization, and Story Generation. Table 9 shows the details of these benchmarks. Additionally, we test our approach on the more challenging open-ended instruction-following benchmarks (MT-Bench, Alpacafarm, and LLMBar), which primarily rely on pairwise comparison task. Unlike the NLG benchmarks, these preference datasets contain more general instructions covering a broader range of tasks with more diverse responses.
Table 1: Kendall ($\tau$) and Spearman ($\rho$) correlation results comparing reference-free, reference-based, and RevisEval methods across natural language generation tasks. This table demonstrates that, without human-annotated references, our proposed RevisEval significantly outperforms reference-free and reference-based methods involving both open-source and proprietary LLM-as-a-Judge.
@@ 4.2 Enhancing LLM-as-a-Judge Performance
For the powerful proprietary LLMs, RevisEval outperforms reference-free and human-annotated reference-based evaluation across tasks with approximately 0.02 in Kendall correlation on average, as demonstrated in Table 1. Notably, in story generation, high-quality human references hinder evaluation for LLM-as-a-Judge, which decreases by 0.06 compared to the reference-free method in the Kendall correlation. In contrast, RevisEval shows that generated response-adapted reference can still significantly enhance evaluation by about 0.08 in Kendall than human reference-based evaluation. An exception is machine translation, where RevisEval aligns closely with reference-based methods, and we analyze this result exists a consistent rationale about reference relevance in Sec 4.5.
@@
Table 2: Results of LLM-as-a-Judge on instruction-following preference tasks. Our proposed RevisEval significantly enhances the performance of both open-source and proprietary LLM-as-a-Judge across various general evaluation tasks. Here, D.R. denotes Direct Response to instruction.
@@
LLM-as-a-Judge can be biased toward longer, verbose answers (Saito et al., 2023; Dubois et al., 2024) or answers that match a similar format (Huang et al., 2024). LLMBar (Zeng et al., 2024) is a challenging benchmark to meta-evaluate such superficial quality biases. As demonstrated in Table 2, these baselines, even after tuning $>100K$ samples, struggle to exceed 50% accuracy, exposing bias challenge significantly. By using response-adapted references from RevisEval, the weak open-source LLM-as-a-Judge’s performance improves substantially by about 6%, showing RevisEval can address superficial quality bias. On proprietary LLM, RevisEval achieves a 3.2% improvement in accuracy compared to reference-free evaluation. Our proposed response-adapted references perform slightly worse than direct responses from the same proprietary model. This is likely because LLMBar emphasizes “instruction-following precision,” where a single ground-truth response exists for the instruction.
@@
In summary, RevisEval consistently outperforms traditional reference-free and reference-based methods, and our revision significantly provides guidance for LLM evaluation by significantly utilizing the generative advantages of LLM.
4.3 Activating classic metrics Performance
Figure 3: Comparative analysis of reference-based metrics performance using references generated by Human/GPT-4 and RevisEval on NLG and instruction following benchmarks. RevisEval can greatly enhance traditional reference-based metrics, even n-gram metrics, reaching performance levels comparable to GPT-4-as-a-Judge in reference-free evaluations.
Since response-adapted references have enhanced LLM-as-a-Judge in Sec. 4.2, this motivates us to ask could these references directly improve the performance of classic metrics, even n-gram metrics?
As the classic text metrics are not compatible with the pairwise comparison, we extend them as follows:
$s=\mathbb{I}(\mathcal{F}_{M}(y_{1},r)>\mathcal{F}_{M}(y_{2},r)),$
@@ 4.3 Activating classic metrics Performance
where $\mathbb{I}(\cdot)$ is the indicator function, determining which $y_{1}$ or $y_{2}$ is preferred based on the metric scores. We use human references in NLG tasks and GPT-4 direct responses in instruction-following benchmarks as the baseline references $\hat{r}$, comparing them with references $r^{\star}$ generated by RevisEval. As shown in Figure 3, RevisEval enables each classic metric to significantly surpass its performance based on baseline references across all tasks. This increase is especially evident in more complex open-ended tasks like story generation and AlpacaFarm. Additionally, we find that using $r^{\star}$ generated by GPT-4 as-a-Reviser combined with classic metrics can yield comparable evaluation performance to GPT-4’s reference-free evaluation. Notably, in AlpacaFarm, when using the response-adapted references, MOVERScore achieves 1.2% improvement compared to the reference-free GPT-4-as-a-Judge.
@@ 4.4 Potential Evaluation Paradigm for Weak Large Language Models
We observe that 1) in the Sec. 4.2, all weak LLMs still exhibit a significant gap compared to GPT-4-as-a-Judge, even after extensive training with high-quality, evaluation-specific data, and 2) in the Sec. 4.3, classic metrics combined with response-adapted references generated by GPT-4 can achieve performance close to the reference-free GPT-4-as-a-Judge. Thus, should we consider a potential evaluation paradigm of “weak LLM-as-a-Reviser + metrics” instead of “weak LLMs-as-a-Judge”?
To explore this, we compare “Llama-as-a-Judge” with “Llama-as-a-Reviser + metrics,” as shown in Table 3. When using references generated by “Llama-as-a-Reviser”, we find that BLEU, METEOR, MOVERScore, and BARTScore can surpass “Llama-as-a-Judge” on average across 3 tasks. Furthermore, we apply a majority voting across multiple metrics, outperforming “Llama-as-a-Judge” over 1.5% on average. This suggests that instead of continuously training weak LLMs to improve their evaluative discrimination capabilities, leveraging their generation abilities for revision may be more effective. Without extra inference costs, this approach can lead to better evaluation outcomes.

### v2

@@ Abstract
With significant efforts in recent studies, LLM-as-a-Judge has become a cost-effective alternative to human evaluation for assessing text generation quality in a wide range of tasks. However, there still remains a reliability gap between LLM-as-a-Judge and human evaluation. One important reason is the lack of guided oracles in the evaluation process. Motivated by the role of reference pervasively used in classic text evaluation, we introduce RevisEval, a novel text generation evaluation paradigm via the response-adapted references. RevisEval is driven by the key observation that an ideal reference should maintain the necessary relevance to the response to be evaluated. Specifically, RevisEval leverages the text revision capabilities of large language models (LLMs) to adaptively revise the response, then treat the revised text as the reference (response-adapted reference) for the subsequent evaluation. Extensive experiments demonstrate that RevisEval outperforms traditional reference-free and reference-based evaluation paradigms that use LLM-as-a-Judge across NLG tasks and open-ended instruction-following tasks. More importantly, our response-adapted references can further boost the classical text metrics, e.g., BLEU and BERTScore, compared to traditional references and even rival the LLM-as-a-Judge. A detailed analysis is also conducted to confirm RevisEval’s effectiveness in bias reduction, the impact of inference cost, and reference relevance.
@@ 1 Introduction
As the large language model (LLM) already exhibits strong alignment with humans (Gilardi et al., 2023; OpenAI et al., 2024), LLM-as-a-Judge (Chang et al., 2024; Li et al., 2024b; Gao et al., 2024b), aka. LLM-evaluator, has emerged as a viable alternative to human evaluation in assessing text generation quality. Given the task instruction and the corresponding model-generated responses, LLMs are prompted to predict preferences or scores for these responses. Despite considerable efforts have been made, such as chain-of-thought (Zheng et al., 2023), specialized rubrics (Liu et al., 2023), and extensive evaluation-specific training datasets (Li et al., 2024a; Wang et al., 2024c; b), human evaluation remains the gold standard in text quality assessment (Zeng et al., 2024) and LLM-as-a-Judge struggles with particular biases (Huang et al., 2024) and being vulnerable to the misleading context (Dubois et al., 2024; Chen et al., 2024). One important reason is the lack of an oracle to direct the evaluation process. Fortunately, classical text evaluation metrics, like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), offer valuable prompts in mitigating such a gap: given an appropriate reference, i.e., the ground-truth answer to the task, calculating the similarity between the references and the model-generated responses can achieve a satisfactory correlation with human evaluations. Furthermore, several studies highlight the reference could prevent being overly sensitive to semantic deficiency  (Sheng et al., 2024) and overcoming bias (Deutsch et al., 2022) in certain cases.
@@
Motivated by the above findings, we deem that an effective reference should maintain high quality while ensuring relevance to the response, which led us to consider that revising the response adaptively could be a good candidate for this reference (Guo et al., 2024). Therefore, we propose a novel evaluation paradigm Revise-and-Evaluation (RevisEval). Specifically, given the (instruction, response) pair, RevisEval first revise the response using the instruction and evaluation rubric, resulting in the response-adapted reference. RevisEval further leverages this generated response-adapted reference to guide final evaluation (e.g., scoring or pairwise comparison). By revising the original response, we ensure that the generated reference is both high-quality and closely relevant to the original content. The comparison between the original and revised responses offers valuable insights for evaluation. Orthogonal to previous work that only focuses on the discrimination of LLMs, RevisEval stands out by fully utilizing the generative potential by revision.
We conduct comprehensive experiments to validate the effectiveness of our proposed RevisEval. Using both proprietary and open-source LLMs, RevisEval consistently achieves better performance compared to reference-free and reference-based evaluation paradigms in both NLG tasks and open-ended instruction-following tasks. Moreover, we seek to verify the effectiveness of the response-adapted references in the classic metrics, e.g., BERT and BERTScore, showing that each metric exceeds itself by up to 3%-10% accuracy compared to using direct response as references. We then combine LLM-as-a-reviser with multiple classic metrics and find that it outperforms (reference-free) LLM-as-a-Judge by over 1.5% on average when using weak LLMs and is comparable when using GPT-4. Finally, we analyze how our paradigm achieves overall superiority. In reducing verbosity and positional bias, our approach offers clear advantages in adversarially designed LLMBar and swapping position testing. Merely increasing inference cost of reference-free evaluation still lags behind RevisEval, demonstrating our method’s efficiency does not rely on naively accumulating cost. Meanwhile, we validate the relationship between reference relevance and efficiency using response-adapted references.
@@ 2.1 Evaluation of Large Language Models
Instruction-tuned LLMs (Ouyang et al., 2022; Dubey et al., 2024; Team et al., 2023) have revolutionized the field of NLP due to their ability to handle a wide range of language-related tasks. Unlike traditional NLP tasks, such as Machine Translation and Summarization, which can be evaluated using N-gram-based metrics like BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) by comparing responses with reference texts, LLMs excel at open-ended text generation tasks (e.g., story generation and instruction-following generation), where no single reference response exists. Consequently, several studies embrace the potential of LLM-as-a-Judge and shift toward reference-free metrics, advocating the abandonment of conventional reference-based evaluation methods (Sheng et al., 2024; Chen et al., 2023). In this paper, we revisit the significance of reference in LLM evaluation. Furthermore, to address the challenge of the absence of a single standard answer in certain evaluation tasks, we propose leveraging LLMs to generate response-adapted references, thereby improving the performance of both traditional metrics and LLM-as-a-Judge.
@@ 3 Methodology
Figure 2: Illustration of our proposed RevisEval. Given an instance ($x,y,a$), we use RevisEval to assess $y$ in rubric $a$. In RevisEval, (i) reviser generates a response-adapted reference $r^{\star}$ by revising the $y$ to enhance the (ii) following LLM-as-a-Judge, even classic text metrics. Here, represents retained segments during the generating response-adapted reference process.
@@ 4 Experiments
In this section, we first present the comprehensive experimental settings in Sec. 4.1 and evaluate the RevisEval in LLM-as-a-Judge across various tasks in Sec. 4.2; we then verify the effectiveness of classic text evaluation metrics when using response-adapted references in Sec. 4.3; building on above findings, we compare two evaluation paradigms, combining LLM-as-a-reviser with multiple metrics and LLM-as-a-Judge, when using weak LLM in Sec. 4.4; finally, we conduct detailed comparative analysis of RevisEval in Sec. 4.5, such as bias reduction, inference cost and reference relevance.
@@ 4.1 Evaluation Settings
We evaluate our approach on multiple classic NLG benchmarks by measuring the correlation between the evaluators/metrics and human annotations in a scoring rating task. We follow the experimental setting of Jiang et al. (2024a) and select four representative NLG tasks and corresponding benchmarks: Data-to-Text (WebNLG), Machine Translation (WMT-22 (zh-en)), Text Summarization (SummEval), and Story Generation (OpenMEVA), and Table 10 shows the details of these benchmarks. Additionally, we test our approach on the more challenging open-ended instruction-following benchmarks (MT-Bench, Alpacafarm, and LLMBar), which primarily rely on pairwise comparison task. Unlike the NLG benchmarks, these preference benchmarks contain more general instructions covering a broader range of tasks with more diverse responses and use accuracy to measure the evaluation performance. Remarkably, in NLG tasks, “ref-based” evaluation relies on human-annotated references as its foundation. In contrast, open-ended instruction-following tasks lack pre-existing references, so “ref-based” evaluation uses machine-generated responses as references to conduct ablation studies verifying the effectiveness of ours references.
Table 1: Kendall ($\tau$) and Spearman ($\rho$) correlation results comparing reference-free, reference-based, and RevisEval methods across natural language generation tasks. This table demonstrates that, without human-annotated references, our proposed RevisEval substantially outperforms reference-free and reference-based methods involving both open-source and proprietary LLM-as-a-Judge.
@@ 4.2 Enhancing LLM-as-a-Judge Performance
For the powerful proprietary LLMs, RevisEval outperforms reference-free and human-annotated reference-based evaluation across tasks with approximately 0.02 in Kendall correlation on average, as demonstrated in Table 1. Notably, in story generation, high-quality human references hinder evaluation for LLM-as-a-Judge, which decreases by 0.06 compared to the reference-free method in the Kendall correlation. In contrast, RevisEval shows that generated response-adapted reference can still greatly enhance evaluation by about 0.08 in Kendall than human reference-based evaluation. An exception is machine translation, where RevisEval aligns closely with reference-based methods, and we analyze this result exists a consistent rationale about reference relevance in Sec 4.5.
@@
Table 2: Accuracy of LLM-as-a-Judge on instruction-following preference tasks. Our proposed RevisEval considerably enhances the performance of both open-source and proprietary LLM-as-a-Judge across various general evaluation tasks. Here, D.R. denotes Direct Response to instruction.
@@
Figure 3: Comparative analysis of reference-based metrics performance using references generated by Human/GPT-4 and RevisEval on NLG and instruction following benchmarks. RevisEval greatly enhance traditional ref-based metrics, even achieving them comparable to GPT-4-as-a-Judge.
@@
LLM-as-a-Judge can be biased toward longer, verbose answers (Saito et al., 2023; Dubois et al., 2024) or answers that match a similar format (Huang et al., 2024). LLMBar (Zeng et al., 2024) is a challenging benchmark to meta-evaluate such superficial quality biases. As demonstrated in Table 2, these baselines, even after tuning $>100K$ samples, struggle to exceed 50% accuracy, exposing bias challenge. By using response-adapted references from RevisEval, the weak open-source LLM-as-a-Judge’s performance improves substantially by about 6%, showing RevisEval can address superficial quality bias. On proprietary LLM, RevisEval achieves a 3.2% improvement in accuracy compared to reference-free evaluation. Our proposed response-adapted references perform slightly worse than direct responses from the same model. This is likely because LLMBar emphasizes “instruction-following precision,” where a single golden response exists for the instruction.
@@
In summary, RevisEval consistently outperforms traditional ref-free and ref-based methods, and our revision provides guidance for LLM evaluation by utilizing the generative advantages of LLM.
4.3 Investigation about Response-adapted References
Table 3: Mean rating for responses and generated references in LLMBar. The detailed rating is in Table 14. Here, Responses 1 and 2 correspond to the position of the response in each pair, and the rating range is in [1, 5].
Text Mean Rating
Response 1 3.19
Response 2 3.25
References (Ours) 4.75
Revision Quality.
As a post-editing mechanism to enhance text quality, we expect the revision to correct errors within the text, including subtle errors (e.g., grammar), as well as higher-level issues (e.g., instruction following precision and factual/logical correctness). LLMBar is an appropriate testbed for conducting detailed case studies and quality assessments. We directly use LLM to score the quality of the responses to be evaluated and the corresponding response-adapted references, focusing on correctness. Table 3 demonstrates that our response-adapted references effectively improve response accuracies. Additionally, the case studies are presented in Appendix J.1; we test RevisEval on JudgeBench (Tan et al., 2024) in Appendix K, focusing on factual/logical correctness and on RewardBench (Lambert et al., 2024) in Appendix N, for more challenging domains.
References Effectiveness.
@@ 4.3 Investigation about Response-adapted References
Traditional reference-based metrics rely heavily on references, directly validating whether the references are effective. Therefore, we use response-adapted references with these metrics to observe evaluation performance across NLG and instruction-following benchmarks. We extend classic reference-based metrics to support pairwise comparisons by using an indicator function to determine preference between $y_{1}$ and $y_{2}$ based on their respective metric scores. We use human references in NLG tasks and GPT-4 direct responses in instruction-following benchmarks as the baseline references, comparing them with ours. As shown in Figure 3, RevisEval enables each classic metric to substantially surpass its performance based on baseline references across all tasks, particularly in more complex open-ended tasks like story generation and AlpacaFarm. Additionally, using references generated by GPT-4 as-a-Reviser combined with classic metrics can yield comparable evaluation performance to GPT-4’s reference-free evaluation. Furthermore, we test response-adapted references in a multi-reference setting, with results detailed in Appendix L and M.
@@ 4.4 Potential Evaluation Paradigm for Weak Large Language Models
We observe that 1) in the Sec. 4.2, all weak LLMs still exhibit a notable gap compared to GPT-4-as-a-Judge, even after extensive training with high-quality, evaluation-specific data, and 2) in the Sec. 4.3, classic metrics combined with response-adapted references generated by GPT-4 can achieve performance close to the reference-free GPT-4-as-a-Judge. Thus, should we consider a potential evaluation paradigm of “weak LLM-as-a-Reviser + metrics” instead of “weak LLMs-as-a-Judge”?
To explore this, we compare “Llama-as-a-Judge” with “Llama-as-a-Reviser + metrics,” as shown in Table 4. When using references generated by “Llama-as-a-Reviser”, we find that BLEU, METEOR, MOVERScore, and BARTScore can surpass “Llama-as-a-Judge” on average across 3 tasks. Furthermore, we apply a majority voting across multiple metrics, outperforming “Llama-as-a-Judge” over 1.5% on average. This suggests that instead of continuously training weak LLMs to improve their evaluative discrimination capabilities, leveraging their generation abilities for revision may be more effective. Without extra inference costs, this approach can lead to better evaluation outcomes. We offer a general guideline: majority voting is recommended as it provides a stable superior evaluation approach, recognizing that no single metric outperforms others across all benchmarks.